### v2

# ResAD: Normalized Residual Trajectory Modeling for End-to-End Autonomous Driving

Abstract
End-to-end autonomous driving (E2EAD) systems, which learn to predict future trajectories directly from sensor data, are fundamentally challenged by the inherent spatio-temporal imbalance of trajectory data. This imbalance creates a significant optimization burden, causing models to learn spurious correlations instead of robust driving logic, while also prioritizing uncertain, distant predictions, thereby compromising immediate safety. To address these issues, we propose ResAD, a novel Normalized Residual Trajectory Modeling framework. Instead of predicting the future trajectory directly, our approach reframes and simplifies the learning task by predicting the residual deviation from a deterministic inertial reference. This inertial reference serves as a strong physical prior, compelling the model to move beyond simple pattern-matching and instead focus its capacity on learning the necessary, context-driven deviations (e.g., traffic rules, obstacles) from this default, inertially-guided path. To mitigate the optimization imbalance caused by uncertain, long-term horizons, ResAD further incorporates Point-wise Normalization of the predicted residual. This technique re-weights the optimization objective, preventing large-magnitude errors associated with distant, uncertain waypoints from dominating the learning signal. On the NAVSIM v1 and v2 benchmarks, ResAD achieves state-of-the-art results of 88.8 PDMS and 85.5 EPDMS with only two denoising steps, demonstrating that ResAD significantly simplifies the learning task and improves planning performance. The code will be released to facilitate further research.
[…]
Figure 1: The core motivation for ResAD: Addressing the Challenge of Imbalanced Trajectory Data. (a) Visualization of the longitudinal data distribution of dataset trajectories under three modeling strategies. Raw Trajectories exhibit significant mean drift and increasing variance, creating a planning horizon dilemma. Our Trajectory Residual centers the distribution, and the Normalized Residual further stabilizes the variance for a simpler, balanced learning objective. (b) Conceptual comparison. Existing methods learn complex raw trajectories directly, risking reliance on spurious correlations. Our ResAD simplifies the task by learning to predict only the necessary residual deviation from a strong physical prior (the inertial reference), focusing the model on context-aware corrections.
1 Introduction
Conventional autonomous driving systems rely on a modular pipeline of perception, prediction, and planning components [liang2020pnpnet, luo2018fast, chitta2022transfuser, wang2022ltp, wu2022trajectory]. This cascaded design is prone to error propagation, leading to suboptimal or unsafe driving. In response to these limitations, End-to-End Autonomous Driving (E2EAD) has emerged as a compelling alternative [hu2023uniad, chen2024end, li2024ego, sima2024drivelm]. E2EAD reframes the driving problem by learning a direct mapping from raw sensor inputs to a future driving trajectory, from which control commands are derived, all within a single, unified framework [jia2023thinktwice, weng2024para, zhou2025opendrivevla, hamdan2025eta].
Recent years have witnessed extensive research into E2EAD methods, focusing on developing more effective representations [jiang2023vad, chen2024vadv2, sun2025sparsedrive, zhang2025bridging], enhancing sensor fusion techniques [Jaeger2023ICCV, ye2023fusionad, chen2024dualat, liu2025gaussianfusion], and designing advanced architectures [yang2024visual, jia2025drivetransformer, song2025don, RAD, zhengdiffusion]. However, these existing methods are all attempting to answer the same question: “What is the future trajectory?” We argue this approach is inherently limited. As shown in Fig.1 (a), the raw trajectory data exhibits a spatio-temporal non-uniformity, which leads to two critical issues that hinder real-world reliability and safety: (1) Spurious Correlations. The immense burden of mapping high-dimensional sensor data directly to a trajectory may cause the model to find shortcuts, relying on spurious correlations instead of the underlying robust logic governing safe driving [jia2023driveadapter, li2024exploring, shao2023reasonnet]. For instance, a model might learn to associate braking with a lead vehicle’s brake lights but fail to understand the red light causing that vehicle to stop, leading it to dangerously follow the car through an intersection. (2) Planning Horizon Dilemma. Trajectory data becomes more uncertain over longer horizons. Consequently, predictions for these distant waypoints often diverge significantly from the eventual ground truth, leading to large loss values during training [salzmann2020trajectron++, veer2022receding, chen2024end, madjid2025trajectory, jiang2025survey]. This skews the optimization process, forcing the model to prioritize large, unpredictable long-term errors over the precision of the critical near-term path essential for immediate collision avoidance.
To address these challenges, we propose  ResAD, a Normalized Residual Trajectory Modeling framework for End-to-End Autonomous Driving. As shown in Fig. 1 (b), our core idea is to decompose the complex prediction task into two distinct components: (1) a deterministic physics-based baseline, the inertial reference, representing the vehicle’s default path in the absence of active control; and (2) a learned residual, the necessary deviations from the inertial reference. By focusing specifically on deviations rather than the entire trajectory, ResAD reframes the learning objective from “What is the future trajectory?” to “Why must the trajectory change?”.This shift significantly simplifies the learning problem. By providing a strong physical prior, the model can focus its capacity on learning the mapping from sensor inputs (e.g., traffic rules, obstacles) to the required corrections (i.e., residuals), rather than expending it on learning complex spatio-temporal dynamics from scratch. To further mitigate the adverse effects of spatial scale variations during optimization, we introduce Point-wise Residual Normalization for the residuals. This technique prevents high-magnitude residuals at certain trajectory points from dominating the learning signal, ensuring that numerically small yet critically important adjustments are properly captured. Additionally, we strategically perturb the ego-vehicle’s state, generating diverse inertial references to counteract planning errors arising from sensor inaccuracies and guide the model toward a broader spectrum of high-quality trajectories. In summary, our contributions are as follows:
  • We revisit the future-trajectory-prediction paradigm in E2EAD, and contend that the spatio-temporal non-uniformity of raw trajectory data leads to spurious correlations and the planning horizon dilemma.
  • We propose ResAD, an E2EAD framework based on the Normalized Residual Trajectory Modeling. It simplifies the learning task by first obtaining an inertial reference and then learning to predict the necessary residual deviations relative to it. We further introduce Point-wise Residual Normalization to mitigate optimization imbalance caused by long-horizon uncertainties.
  • Extensive experiments and analyses validate the effectiveness of the proposed ResAD. On the NAVSIM v1 and v2 benchmark, our method achieves state-of-the-art performance with scores of 88.8 for PDMS and 85.5 for EPDMS.
2 Related Work
2.1 End-to-End Autonomous Driving
End-to-end autonomous driving (E2EAD) seeks to overcome the limitations of traditional modular pipelines, such as error accumulation and inter-module information loss [jia2023driveadapter, min2024driveworld, hanselmann2022king, nie2024reason2drive, jia2024bench2drive, song2025don, sun2025generalizing]. Pioneering works like UniAD [hu2023uniad] introduced a planning-oriented architecture that jointly optimizes perception and forecasting to mitigate error propagation. VAD [jiang2023vad] further streamlined the pipeline with a fully vectorized scene representation, enabling the enforcement of explicit, instance-level safety constraints. More recently, generative models have become a new frontier in E2EAD research [chi2023diffusion, chen2024ppad, feng2025artemis, zheng2024genad, liao2025diffusiondrive, xu2025knowledge, jiang2025diffvla, li2025discrete]. GoalFlow [xing2025goalflow] introduces a goal-conditioned generative model that first selects an optimal goal point based on scene context and then uses Flow Matching to efficiently generate high-quality trajectories towards it. Despite these advances, existing methods predominantly rely on the direct prediction of future trajectories. In this work, we depart from this paradigm by introducing Normalized Residual Trajectory Modeling. Our method formulates a trajectory by decomposing it into a physics-based inertial reference and a learnable residual, offering a more structured and interpretable approach to trajectory representation.
Figure 2: The proposed ResAD framework. Instead of predicting the entire trajectory, ResAD establishes a strong physical prior based on the vehicle’s current state: the Inertial Reference. By applying Inertial Reference Perturbation, the framework generates a diverse set of initial intent hypotheses. Finally, the diffusion decoder (DiffDecoder), conditioned on these references, learns to predict the necessary Normalized Residuals. We highlight the 1st-ranked and 5th-ranked output trajectories, denote as Top-1 Traj and Top-5 Traj.
2.2 Multimodal Planning
Most E2EAD systems produce a single, deterministic trajectory, an approach that struggles with the inherent diversity of real-world driving scenarios. To address this, several works have explored multimodal planning [chen2024vadv2, li2025generalized, song2025breaking, guo2025ipad]. VADv2 [chen2024vadv2] proposes a probabilistic planning framework that outputs a distribution of future trajectories, which can be sampled to produce a diverse set of behaviors. The Hydra-MDP series [li2024hydra, li2025hydra++] employs policy distillation to select multiple trajectories from a vocabulary guided by an expert. GTRS [li2025generalizedtrajectoryscoringendtoend] adopts a different strategy, scoring a set of pre-generated trajectories to ensure both diversity and safety. DiffusionDrive [liao2025diffusiondrive] highlights the challenge of mode collapse in generative-based models, addressing it by anchoring trajectory generation to a fixed cluster vocabulary. However, these methods fundamentally rely on a static, predefined vocabulary. This makes them both inefficient and restrictive, forcing them to evaluate irrelevant options while being unable to generate truly optimal trajectories outside the discrete set. In contrast,  ResAD benefits from its unique trajectory modeling strategy, which enables it to directly denoise from the Gaussian noise. This approach yields superior, context-aware multimodal trajectories without relying on a fixed vocabulary.
3 Methodology
3.1 Preliminaries
E2EAD aims to learn a unified policy, $\pi$, that directly maps raw sensor inputs, $\mathcal{O}$, to a sequence of waypoints, $\tau=\{(x_{t},y_{t})\}_{t=1}^{T_{f}}$, where $T_{f}$ denotes the planning horizon, and $(x_{t},y_{t})$ is the predicted future location of each waypoint at time $t$. We construct the proposed ResAD using a vanilla diffusion [ho2020denoising, songdenoising] framework. The diffusion model defines a Markovian chain of diffusion forward process $q$ by gradually adding noise to sample data $z_{0}$, over a series of $T$ timesteps, which can be formulated as:
$q(\bm{z}_{t}|\bm{z}_{0})=\mathcal{N}(\bm{z}_{t}|\sqrt{\bar{\alpha}_{t}}\bm{z}_{0},(1-\bar{\alpha}_{t})\bm{I}),$ (1)
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}.$ The hyperparameters $\beta_{t}$ controlling the amount of noise added at each step. As $t\xrightarrow{}T$, $z_{T}$ approaches a pure Gaussian noise distribution. We train the denoise model $\pi_{\theta}$ to predict $z_{0}$ from $z_{i}$ with the guidance of conditional information $c$, where $\theta$ are the trainable parameters. At the inference stage, the trained network is used to iteratively denoise a pure noise $z_{T}\sim\mathcal{N}(0,\mathbf{I})$ to produce a clean data sample $z_{0}$, which is defined as:
$p_{\theta}\left(z_{0}\mid c\right)=\int p\left(z_{T}\right)\prod_{i=1}^{T}p_{\theta}\left(z_{i-1}\mid z_{i},c\right)\mathrm{d}z_{1:T}.$ (2)
In this work, instead of directly generating the future trajectory points, we define the data samples as a set of normalized residuals.
3.2 Normalized Residual Trajectory Modeling
As illustrated in Fig. 2, ResAD takes multi-view images and LiDAR point clouds as input, which are fused by a Transfuser-style encoder [chitta2022transfuser]. We first generate an inertial reference from the ego-vehicle’s current state.  ResAD then perturbs this reference to generate a cluster, ensuring robustness and enabling multimodal predictions. Finally, the Diffusion Decoders employ cross-attention to merge the encoded features, using the inertial reference cluster as a condition to guide the denoising process.
Trajectory Residual Modeling. The core idea of ResAD is to reframe the complex trajectory prediction task into a simpler, more interpretable learning problem. Instead of predicting the entire future trajectory from scratch, we use a constant velocity mode [rhinehart2018r2p2] to extrapolate the inertial reference trajectory from the ego-vehicle’s current state, and predict the residual. This remodeling compels the model to learn the control interventions required to deviate from the default reference, focusing on the critical, high-level driving decisions needed by the scene.
Let the ego-vehicle’s velocity be $\bm{v}_{0}=(v_{x,0},v_{y,0})$ and its position be $\mathbf{p}_{0}=(x_{0},y_{0})$ at the current time $t=0$. The inertial reference trajectory $\tau_{\mathrm{ref}}$ for future timesteps $t_{i}$ in the prediction horizon $T_{f}=\{t_{1},t_{2},...,t_{N}\}$ is calculated as:
$\mathbf{p}_{t_{i}}=\mathbf{p}_{0}+\mathbf{v}_{0}\cdot\Delta t_{i}.$ (3)
This reference $\tau_{\mathrm{ref}}$ represents the path the vehicle would follow with no control inputs. We define the trajectory residual $\bm{r}$ as the point-wise difference between the ground-truth trajectory $\tau_{\mathrm{gt}}$ and the reference trajectory $\tau_{\mathrm{ref}}$:
$\bm{r}=\tau_{\mathrm{gt}}-\tau_{\mathrm{ref}}.$ (4)
This residual $\bm{r}$ quantifies the precise corrections a human driver applied to navigate the environment. The learning objective of ResAD is thus to predict this residual, effectively capturing the driver’s decision-making process.
Point-wise Residual Normalization. A key challenge in planning is the scale variance of coordinates across the time horizon. Points further in the future have numerically larger values, which can cause the optimization to be dominated by far-field errors, neglecting the fine-grained, safety-critical adjustments required in the near field. As shown in Fig. 1(a), while residual modeling mitigates this by focusing on deviations, the scale issue within the residuals themselves persists. We propose Point-wise Residual Normalization (PRNorm) to resolve this.
Given a residual trajectory $\bm{r}$, which is a sequence of $T_{f}$ displacement vectors $\{r_{1},r_{2},\dots,r_{T_{f}}\}$, where each $r_{t}=(r_{t}^{x},r_{t}^{y})$ is a 2D vector. A standard min-max scaling is performed on a component-wise basis for each dimension $d\in\{x,y\}$. The extremal values, $r_{\min}^{d}$ and $r_{\max}^{d}$ are pre-computed across all timesteps and all trajectories in the entire training dataset:
$r_{\min}^{d}=\min_{j,t}(r_{j,t}^{d}),\quad r_{\max}^{d}=\max_{j,t}(r_{j,t}^{d}),$ (5)
where $j$ indexes trajectories in the training set and $t$ indexes the timestep. These values define the tightest axis-aligned bounding box for the residual. To provide fine-grained control over the final feature distribution, we introduce a hyperparameter $\gamma>0$. This parameter defines the bounds of the symmetric output interval $\gamma,\gamma]$. The complete transformation of PRNorm for each component $r_{t}^{d}$ of every vector $r_{t}$ is given by:
$\tilde{r}_{t}^{d}=2\gamma\left(\frac{r_{t}^{d}-r_{\mathrm{min}}^{d}}{r_{\mathrm{max}}^{d}-r_{\mathrm{min}}^{d}+\epsilon_{0}}\right)-\gamma.$ (6)
The small constant $\epsilon_{0}$ is added to the denominator to ensure numerical stability. Through this, we can get the normalized residual $\tilde{\bm{r}}=\text{PRNorm}(\bm{r})$.
Table 1: Performance on the NAVSIM v1 navtest Benchmark.
[…]
“C”: Camera, “L”: LiDAR. The best and second-best scores are highlighted in bold and underlined, respectively.
Inertial Reference Perturbation. Driving is an inherently multi-modal task. Most methods depend on a fixed trajectory vocabulary, where most options are irrelevant to the current scene, causing inefficiency. ResAD generates multimodal trajectories by perturbing Inertial Reference. This approach offers a dual benefit. First, it generates diverse intent hypotheses by creating varied inertial references and corresponding residuals. Second, it compels the model to handle noise inherent in ego-sensors such as GPS and IMU.
Specifically, we introduce stochastic perturbations directly into the initial velocity $\bm{v}_{0}$. We generate $K$ distinct perturbation vectors $\bm{\delta}_{\mathbf{v},k}$ by sampling from a zero-mean multivariate Gaussian distribution:
$\bm{\delta}_{\mathbf{v},k}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma})\quad\text{for}\quad k=1,\ldots,K.$ (7)
Here, the covariance matrix $\bm{\Sigma}=\mathrm{diag}(\sigma_{vx}^{2},\sigma_{vy}^{2})$ governs the variance of the perturbations along the longitudinal and lateral axes, respectively. These hyperparameters effectively define the exploration scope of our model’s initial hypotheses. Each perturbation is additively fused with the original velocity vector to forge $K$ unique, perturbed initial states. By propagating each of these perturbed velocity vectors $\mathbf{v}_{0,k}^{\prime}$ through the constant velocity model Eq. 3, we generate a set of $K$ distinct inertial references $\{\tau_{\mathrm{ref},~k}\}_{k=1}^{K}$ and corresponding residuals:
$\mathbf{v}_{0,k}^{\prime}$ $=\mathbf{v}_{0}+\mathbf{\delta}_{\mathbf{v},k},$ (8)
$\{\bm{r}_{k}\}_{k=1}^{K}$ $=\{\tau_{\mathrm{gt}}-\tau_{\mathrm{ref},~k}\}_{k=1}^{K}.$
Training and Inference. In training, adding Gaussian noise to the residual cluster normalized by PRNorm:
$\bm{z}{{}_{k}^{(i)}}=\sqrt{\bar{\alpha}_{i}}\tilde{\bm{r}}_{k}+\sqrt{1-\bar{\alpha}_{i}}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),$ (9)
where $\tilde{\bm{r}}_{k}=\mathrm{PRNorm}(\bm{r}_{k})$. The diffusion decoder $f_{\theta}$ takes $K$ noisy normed trajectory residuals to generate denoised residuals $\{\hat{\bm{r}}_{k}\}_{k=1}^{K}$:
$\{\hat{\bm{r}}_{k}\}_{k=1}^{K}=f_{\theta}(\{\bm{z}^{(i)}_{k}\}_{k=1}^{K},c),$ (10)
where $c$ represents the conditional information. Note that $c$ is composed of query features extracted from the encoder and the corresponding timestep embedding. Crucially, $c$ also incorporates unique positional encoding features derived from each of the perturbed inertial references. These encodings are essential for the model to distinguish between the different intent hypotheses and subsequently generate a diverse set of trajectories. The diffusion loss is computed as:
$\mathcal{L}_{\mathrm{diff}}=\sum_{k=1}^{K}\mathcal{L_{\mathrm{rec}}}(\hat{\bm{r}}_{k},\bm{r}_{k}),$ (11)
here, $\mathcal{L}_{\mathrm{rec}}$ can be a simple L1 loss or MSE loss.
During inference, the denoising process starts with $K_{\mathrm{infer}}$ Gaussian noise to generate residuals. We take 2 timesteps to get the final predictions $\{\hat{\bm{r}}_{k}\}_{k=1}^{K_{\mathrm{infer}}}$ by DDIM [songdenoising]. Then the predicted residuals are added to the corresponding perturbed inertial reference to get the multi-modal trajectory $\{\hat{\tau}_{k}\}_{k=1}^{K_{\mathrm{infer}}}$.
3.3 Multimodal Trajectory Ranker
Inspired by VADv2 [chen2024vadv2] and Hydra-MDP [li2024hydra], we develop a Trajectory Ranker to select the optimal trajectory from multiple candidates generated by the planning module. Given a set of trajectory candidates $v_{k}$, where k is the number of trajectories, we feed them into a Transformer to facilitate interaction with the perception representations, $E_{env}$, which can be expressed as follows:
$\mathcal{V}$ $=\mathrm{PosEmb}(v_{k}),$ (12)
$\mathcal{V}^{\prime}$ $=\mathrm{Transformer}(Q=\mathcal{V},K,V=E_{env})+E.$
$\mathrm{PosEmb}(\cdot)$ denotes the position embedding, and ego status $E$ is embedded into the transformer output. Subsequently, the latent vector $\mathcal{V}^{\prime}$ is fed into a set of MLP heads to predict the score $\{\hat{\mathcal{S}}_{i}^{m}|i=1,...,k\}_{m=1}^{|M|}$ for each metric $m\in M$ and the $i$-th trajectory, where M represents the set of metrics used in PDMS or EPDMS. The ranker is trained with the ground truth score $\{{\mathcal{S}}_{i}^{m}|i=1,...,k\}_{m=1}^{|M|}$ to distill the knowledge from the rule-based planner and the ground truth waypoints as follows:
$\mathcal{L}_{\mathrm{ranker}}=\sum_{i=1}^{k}y_{i}\log(\hat{\mathcal{S}}_{i}^{im})+\sum_{m,i}\text{BCE}({S}_{i}^{m},\hat{\mathcal{S}}_{i}^{m}),$ (13)
where $y_{i}=\frac{e^{-({\tau}_{\mathrm{gt}}-\hat{\tau}_{i})^{2}}}{\sum_{j=1}^{k}e^{-({\tau}_{\mathrm{gt}}-\hat{\tau}_{j})^{2}}}.$ During inference, we compute scores for the outputs of the planning head and select the trajectory with the highest weighted score as the final output.
4 Experiments
Table 2: Performance on the NAVSIM v2 navtest Benchmark with Extended Metrics.
[…]
4.1 Benchmark
We evaluate the proposed ResAD on the NAVSIM v1 [dauner2024navsim] and NAVSIM v2 [Cao2025ARXIV] benchmark. NAVSIM is built upon the real-world NuPlan dataset [karnchanachari2024towards] and exclusively features relevant annotations and sensor data sampled at 2 Hz. The NAVSIM dataset contains two parts: navtrain and navtest, including 1192 and 136 scenarios respectively, used for trainval and test.
NAVSIM v1. In this benchmark, each predicted trajectory is sent to a simulator, which validates the driving metrics in the corresponding environment. The planning capabilities of models are assessed using the PDM score (PDMS), which is calculated as follows:
$\mathrm{PDMS}=\mathrm{NC}\times\mathrm{DAC}\times\frac{(5\times\mathrm{TTC}+2\times\mathrm{C}+5\times\mathrm{EP})}{12},$ (14)
where the sub-metrics NC, DAC, TTC, C, EP represent the No At-Fault Collisions, Drivable Area Compliance, Time to Collision, Comfort, and Ego Progress.
NAVSIM v2. A new Extended PDM Score (EPDMS) is introduced in NAVSIM v2, which can be formulated as:
$\mathrm{EPDMS}=\mathrm{NC}\times\mathrm{DAC}\times\mathrm{DDC}\times\mathrm{TL}\times\\ \frac{(5\times\mathrm{TTC}+2\times\mathrm{C}+5\times\mathrm{EP}+5\times\mathrm{LK}+5\times\mathrm{EC})}{22}.$ (15)
The extended sub-metrics DDC, TL, LK, and EC correspond to the Driving Direction Compliance, Traffic Lights Compliance, Lane Keeping Ability, and Extended Comfort.
4.2 Implementation Details
For fair comparison, our model adopts an identical perception module and ResNet-34 [resnet] or V2-99 [v299] as its backbone following existing methods [chitta2022transfuser, li2025hydra++]. The proposed ResAD is equipped with 2 cascaded diffusion decoder layers. We set the number of modes $K_{\mathrm{train}}=20$ for training and $K_{\mathrm{infer}}=200$ for testing. The model is trained from scratch on the navtrain split for 100 epochs using the DDPM [ho2020denoising], with a timestep $T$ of 1000. The training is distributed across 8 NVIDIA L20 GPUs, with a total batch size of 512, and is optimized using AdamW. In inference, we use DDIM [songdenoising] to sample the predictions with only 2 denoising steps. The resulting candidates are then evaluated by the trained ranker, which selects the highest-scoring trajectory as the output. We predict $T_{f}=8$ timesteps and the interval between each time step is 0.5s. For more details, please refer to the supplementary material.
Figure 3: Visual comparison of ResAD. This figure compares the 20 trajectory candidates from ResAD and DiffusionDrive. DiffusionDrive relies on a static, context-agnostic vocabulary, often proposing infeasible or irrelevant trajectories (highlighted by red circles). In contrast, ResAD dynamically generates a set of context-aware trajectories via IRP. This demonstrates our method’s more efficient multimodal exploration, which avoids wasting capacity on invalid options and leads to a higher multimodal planning quality ($\mathcal{P}_{m}$), as validated in Sec. $4.$
4.3 Main Results
Quantitative Comparison. The results presented in Tab. 1 show that ResAD achieves a SOTA performance on the NAVSIM v1 navtest split, with a PDMS of 88.8 using a ResNet-34 backbone, surpassing strong baselines such as DiffusionDrive (PDMS 88.1) and WoTE (PDMS 88.3) under the same backbone. With a stronger V2-99 backbone, ResAD further improves to PDMS 90.6, surpassing GoalFlow and Hydra-MDP. On the more challenging NAVSIM v2 benchmark, the advantages of ResAD are further extended. As shown in Tab. 2ResAD reaches the best or second-best performance across almost all extended sub-metrics. Relative to DiffusionDrive,  ResAD improves notably on EP (88.2 vs. 87.5) and DAC (97.2 vs. 95.9), indicating stronger route completion and adherence to drivable areas under the richer EPDMS protocol.
Qualitative Comparison. A qualitative comparison on NAVSIM (Fig. 3) highlights the different multimodal strategies of ResAD and DiffusionDrive. While both successfully avoid the mode collapse typical of vanilla diffusion, their underlying approaches diverge significantly. DiffusionDrive relies on a static, predefined trajectory vocabulary. This context-agnostic approach forces it to generate many irrelevant or unfeasible options, such as proceeding straight in a sharp turn scenario (highlighted by red circles in the figure). Although a subsequent filtering step can prune these invalid paths, this two-stage process is inherently inefficient. In contrast, ResAD overcomes this limitation through its distinct trajectory modeling shift. This is achieved through a mechanism of perturbing the ego-vehicle’s velocity. It directly explores a set of plausible behaviors, generating trajectories that are inherently consistent with the immediate driving context. A real-world vehicle demonstration of the proposed method is available in the supplementary material.
4.4 Ablation Studies and Analysis
Table 3: Ablation study on the influence of each component.
Model Description $\textbf{NC}\uparrow$ $\textbf{DAC}\uparrow$ $\textbf{EP}\uparrow$ $\textbf{TTC}\uparrow$ $\textbf{C}\uparrow$ $\textbf{PDMS}\uparrow$
$\mathcal{M}_{0}$ Base Model 97.8 94.2 78.1 93.4 100 84.9
$\mathcal{M}_{1}$ $\mathcal{M}_{0}$ + Ranker 98.3 94.3 77.8 94.6 100 85.1
$\mathcal{M}_{2}$ $\mathcal{M}_{1}$ + TRM 97.4 96.6 80.3 93.2 100 86.3
$\mathcal{M}_{3}$ $\mathcal{M}_{2}$ + PRNorm 97.6 96.7 81.4 93.3 100 87.2
$\mathcal{M}_{4}$ $\mathcal{M}_{3}$ + IRP 98.0 97.5 83.4 94.1 100 88.8
Figure 4: PRNorm stabilizes optimization and accelerates convergence. This figure compares the training dynamics of ResAD with and without PRNorm.
Component Analysis. As summarized in Tab. 3, adding the Ranker to the baseline ($\mathcal{M}_{0}\rightarrow\mathcal{M}_{1}$) yields only marginal benefit (PDMS +0.2) and mildly improves safety, indicating its choices concentrate in a narrow mode. Introducing Trajectory Residual Modeling (TRM) delivers the key boost (DAC +2.3, EP +2.5, PDMS +1.2). By providing a strong physical prior, TRM compels the model to learn the underlying driving logic rather than relying on the spurious correlations it might learn from raw trajectory data. This is further qualitatively validated in Fig. 5. Adding Point-wise Residual Normalization (PRNorm) further refines performance and stabilizes convergence, improving EP to 81.4 and PDMS to 87.2. PRNorm stabilizes optimization by resolving spatio-temporal scale variance, preventing far-field errors from dominating. This refines the trajectory quality, leading to another notable increase in EP (+1.1). Finally, Inertial Reference Perturbation (IRP) delivers the largest single improvement, reaching 88.8 PDMS (+1.6 over $\mathcal{M}_{3}$). IRP unlocks true, context-aware multimodality. It generates a diverse yet plausible set of intent hypotheses, providing high-quality candidates for the ranker to evaluate. This is the key that enables the model to excel in complex scenarios, dramatically improving EP (+2.0) and DAC (+0.8).
Impact of PRNorm on Training Dynamics. As shown in Fig. 4, PRNorm enables a significantly faster decline in loss compared to the baseline (vanilla min-max normalization), accelerating model convergence. Furthermore, we calculate the PDMS of the predicted trajectory every step, which is also higher with PRNorm. It demonstrates its comprehensive benefits to both training efficiency and the final performance.
Qualitative Analysis of the TRM. As shown in Fig. 5, we remodel the trajectory as a residual against the inertial reference (IR). The IR can serve as a counterfactual baseline: in steady cruising, predictions largely coincide with the IR, yielding small residuals and stable optimization; in dynamic scenes (merging, intersections, sharp turns, braking), the baseline becomes highly inaccurate, producing strong, directional residuals. These residual vectors point away from obstacles or follow road curvature, indicating the model learns why to deviate (due to obstacles, yielding, or turning) rather than merely fitting a complex trajectory. Moreover, IRP samples a bundle of lightly perturbed baselines around the original IR, expanding the physically plausible neighborhood. The model learns sharper residuals over these references, avoiding mode collapse. Together, IR + IRP focus capacity on what must change and encourage exploration of a more effective trajectory space, enabling robust predictions with few denoising steps.
Figure 5: The impact of Inertial Reference. We highlight the 1st-ranked and 5th-ranked trajectory with their corresponding IRs. Red arrows are used to highlight key examples of the predicted residual.
Table 4: Extending Normalized Residual Trajectory Modeling to other models.
Model $\textbf{NC}\uparrow$ $\textbf{DAC}\uparrow$ $\textbf{EP}\uparrow$ $\textbf{TTC}\uparrow$ $\textbf{C}\uparrow$ $\textbf{PDMS}\uparrow$
Transfuser 97.7 92.8 79.2 92.8 100 84.0
[…]
TransfuserDP 97.4 93.5 79.0 93.0 100 84.5
[…]
Table 5: Comparison of running time and performance. FPS and total runtime are measured on an NVIDIA 4090 GPU.
[…]
TransfuserDP 84.6 20 84.4 130.0ms 101M 7
DiffusionDrive 88.1 20 60.3 7.6ms 60M 45
ResAD w/o Ranker - 20 86.1 7.7ms 62M 45
ResAD 88.3 20 86.1 11.4ms 68M 37
Superiority of the Normalized Residual Trajectory Modeling. We demonstrate the numerous advantages of Normalized Residual Trajectory Modeling (NRTM) through further experiments: (1) Plug-and-Play. NRTM is a drop-in strategy that requires limited changes to the architecture or loss design: attach TRM (+PRNorm) to any planner and train as usual. As shown in Tab. 4, it consistently lifts performance on both an MLP planner (Transfuser) and a diffusion planner (TransfuserDP). On Transfuser, TRM raises PDMS from 84.0 to 85.2, and TRM+PRNorm further to 85.6; comparable gains are observed on TransfuserDP. By residual modeling, NRTM captures essential high-order driving primitives and, via spatio-temporal data alignment, accelerates optimization. It delivers method-agnostic improvements to E2EAD safety and efficiency with minimal integration cost. (2) Efficient and Context-aware Multimodal Planning. As shown in Tab. 5, compared with DiffusionDrive, all increases in parameters and inference time for ResAD come from the Ranker, which selects the optimal trajectory from a set of candidates. Beyond reporting PDMS, we also introduce the mean PDMS $\mathcal{P}_{m}$ to evaluate the average quality of the model’s multimodal trajectory outputs. ResAD attains a substantially higher $\mathcal{P}_{m}$, providing further quantitative evidence that ResAD generates context-aware multimodal trajectories. (3) Test-Time Scaling.
Figure 6: Dynamic number of ${K}_{\text{infer}}$. This figure plots PDMS as a function of the inference-time sampled noise numbers ${K}_{\text{infer}}$.
As shown in Fig. 6, while DiffusionDrive exhibits a distinct performance drop when $K_{\text{infer}}<K_{\text{train}}$ and remains around 88.1 when $K_{\text{infer}}>20$,  ResAD realizes a consistent performance and continuous growth. The introduction of IR provides effective physical prior, constraining predictions within a context-relevant region and effectively avoiding invalid samples. Therefore, NRTM not only boosts the performance but also converts extra inference compute into consistent PDMS gains, revealing a compute-adaptive, anytime advantage at deployment.
5 Conclusion
In this work, we revisit the conventional E2EAD paradigm, highlighting a key challenge where directly predicting raw trajectories can create a severe optimization imbalance and risks learning spurious correlations. We propose ResAD, a novel framework that fundamentally reframes the learning problem. Our approach introduces two key innovations: Trajectory Residual Modeling (TRM) and Point-wise Residual Normalization (PRNorm). TRM provides a strong physical prior via an inertial reference, forcing the model to learn only the necessary, context-driven residuals. This simplifies the task and focuses capacity on learning robust driving logic, not spurious patterns. PRNorm mitigates the optimization imbalance, ensuring that safety-critical, near-term adjustments are prioritized over uncertain long-term predictions. Our state-of-the-art results on NAVSIM are a direct consequence of this conceptual shift. By taming the imbalanced data and simplifying the learning objective, ResAD provides a significantly more robust, stable, and scalable foundation for E2EAD systems.